Write the expression for the channel capacity C of a communication system. If S/N =15 and
B =10,000, calculate C.
1
Expert's answer
2019-06-03T05:51:37-0400
"C=B \times log2(1+S/N)=10 000 \times log2(1+15)=10 000 \times 4=40 000 bit/sec"